The sum of the digits of a two-digit number is 7. The original two-digit number is 3 less than 4 times the number with its digits reversed. Find the original two-digit number.
: The sum of the digits of a two-digit number is 7. The original two-digit number is 3 less than 4 times the number with its digits reversed. Find the original two-digit number.

Answer by stanbon(17617)

The sum of the digits of a two-digit number is 7.
The original two-digit number is 3 less than 4 times the number with its digits reversed.
Find the original two-digit number.
----------------
Let the original number be 10t+u where t is the tens digit and u is the units digit.
--------
EQUATIONS:
t + u = 7
10t+u = 4(10u+t)-3
------------
Rearrange:
t + u = 7
6t -39u = -3
----------------
Modify:
6t + 6u = 42
6t - 39u = -3
---------------
Subtract to solve for "u":
45u = 45
u = 1
--------
Substitute into t+u = 7 to solve for "t":
t+1 = 7; t=6
------------------
Original Number: 61
======================
Cheers,
Stan H.